# Spatial Scale of Agglomeration and Dispersion Number, Spacing, and the Spatial Extent of Cities

### v7

@@ 2.2   The Braid model: A “global” dispersion force
Next, we consider another minimal model with a “global” dispersion force. We replace the local congestion force in the Beckmann model with a trade-based dispersion force:
$v_{i}(\bm{x})=y_{i}(\bm{x})\Big{(}\sum_{j}\phi_{ij}x_{j}\Big{)}^{\alpha},$ (6)
where $y_{i}(\bm{x})$ is the income of an agent in region $i$ under given $\bm{x}$, which depends on costly transport. Specifically, the economy features gravity-type trade flows, whose microfoundations abstracted away for brevity. Each region $m$ is endowed with one unit of expenditure, which is allocated across all agents according to accessibility. The share received by an agent located in region $i$ from region $m$ is given by
$s_{i\mid m}\equiv\frac{\phi_{im}}{\sum_{l\in\mathcal{I}}x_{l}\phi_{lm}},$ (7)
ensuring that total spending from region $m$ equals one: $\sum_{i\in\mathcal{I}}x_{i}s_{i\mid m}=1$. The income of an agent in region $i$ is then given by
$y_{i}(\bm{x})=\sum_{m\in\mathcal{I}}s_{i\mid m}=\sum_{m\in\mathcal{I}}\frac{\phi_{im}}{\sum_{l\in\mathcal{I}}x_{l}\phi_{lm}}.$ (8)
For convenience, we refer this reduced-form model [^6] The model does not specify how consumers’ utility depends on their consumption of the tradables, where the expenditure comes from, what is the technology of producing tradables, how the prices of tradables are determined in equilibrium. as the Braid model after Braid-RSUE1988 who studied the case $\alpha=0$.
The uniform distribution is an equilibrium. Evaluating Eq. 3 gives
@@
Agglomeration occurs as long as $\alpha>0$. If $10<\alpha<$ 1, $\bar{\bm{x}}$ is stable for high transport costs (large $\Theta$Leftrightarrow small $\phi$) and unstable for low transport costs (small $\Theta$Leftrightarrow large $\phi$). [^7] If $\alpha\geq 1$, the symmetry is unstable for all $\phi\in(0,1)$, as $\Theta<1$ implies $\omega=(\alpha-\Theta)\Theta>0$.
Importantly, $\omega$ features a “global” dispersion force—a negative term in $\Theta$. The dispersion force stems from an interregional crowding effect embedded the share formula (Eq. 7). For an agent in region $i$, an increase in the mass of agents in any location implies less income from any other region $m$ because $s_{i\mid m}$ is smaller if region $m$ has better access to other locations (i.e., if $\sum_{l\in\mathcal{I}}x_{l}\phi_{lm}$ is larger). Concretely, we see
$\frac{\partial s_{i\mid m}}{\partial x_{j}}=-\frac{\phi_{im}}{\sum_{l\in\mathcal{I}}x_{l}\phi_{lm}}\times\frac{\phi_{jm}}{\sum_{l\in\mathcal{I}}x_{l}\phi_{lm}}=-s_{i\mid m}\times s_{j\mid m}<0$ (10)
for all $j$. The magnitude of this negative effect depends on both $\phi_{im}$ and $\phi_{jm}$. As such, this must appear as a second-order term of the transport cost index $\Theta$.
2.3   Contrasting implications of transportation costs
@@ D The Braid model with asymmetry
Figure 3: Equilibrium values of $x_{1}$ in the Braid model and the Beckmann model
@@
Evidently, the Braid model and the Beckmann model offer contrasting implications regarding how transport costs shape endogenous asymmetries in the spatial distribution. Figures 3B and 3A highlight the contrast by showing the full equilibrium paths over the $\phi$-axis. In Fig. 3A, the symmetric equilibrium in the Beckmann model is stable if transport costs are low (i.e., if $\phi$ is large), but agglomeration emerges as transport costs increase (as $\phi$ decreases). The Braid model exhibits the opposite pattern, as shown in Fig. 3B.
These tendencies persist under the presence of asymmetries. In Figs. 3D and 3C, region 1 has an exogenous advantage. The symmetric pattern is not an equilibrium anymore. However, focusing on the equilibrium paths with $x_{1}>x_{2}$, broad implications are the same as the symmetric case. Lowering transport costs leads to dispersion in the Beckmann model, but to agglomeration in the Braid model.
While the literature has not explicitly referred to the spatial nature of dispersion forces, it is recognized that various dispersion mechanisms have opposing effects. In particular, Fujita-Thisse-Book2013, Chapter 8, compared the seminal models by Krugman-JPE1991 and Helpman-Book1998 and noted that “Krugman’s scenario is reversed” (p.289) in Helpman-like frameworks with urban costs. That is, agglomeration is preferred to dispersion if transport costs are sufficiently low in the former, while the converse is true for the latter. In fact, the minimal models capture the essence of these frameworks, as discussed in the remainder of this section.
@@ 2.4   Benefit matrix and the spatial scale of economic forces
The discussed contrast illustrates how the spatial scale of dispersion forces—local or global—can alter the spatial implications of declining transport costs. To formalize this insight and provide a tool for analyzing more sophisticated models, we now introduce the benefit matrix. For each specific indirect utility function $\bm{v}$, we define its benefit matrix as its elasticity matrix at the symmetric equilibrium, $\mathbf{V}\equiv\frac{\bar{x}}{\bar{v}}[\frac{\partial v_{i}}{\partial x_{j}}(\bar{\bm{x}})]_{i\in\mathcal{I},j\in\mathcal{I}}$. It summarizes how changes in the population distribution affect individual utility in the model under consideration. This can be formalized as follows:
Observation 1.
@@
For each specification of $\bm{v}$, the utility gain $\omega$ at the symmetric equilibrium is an eigenvalue of the benefit matrix $\mathbf{V}$, and the corresponding eigenvector is $\bm{z}=(1,-1)$. ∎
[…]
$\mathbf{V}=-\beta\mathbf{I}+\alpha\mathbf{D}\quad\text{with}\quad\mathbf{D}=\frac{1}{1+\phi}\begin{bmatrix}1&\phi\\ \phi&1\end{bmatrix},$ (11)
@@
where $\mathbf{I}$ is the identity matrix and $\mathbf{D}$ is the row-normalized proximity matrix. Observe that the expression for the utility gain, $\omega=-\beta+\alpha\Theta$, mirrors the structure of $\mathbf{V}$. In fact, the index $\Theta=\frac{1-\phi}{1+\phi}$ is an eigenvalue of $\mathbf{D}$, with the eigenvector being $\bm{z}$. It follows that $\mathbf{V}\bm{z}=(-\beta+\alpha\Theta)\bm{z}=\omega\bm{z}$, showing that $\omega$ is indeed an eigenvalue of $\mathbf{V}$.
Importantly, the representation of $\mathbf{V}$ based on $\mathbf{D}$ succinctly summarizes how agglomeration and dispersion forces interact with transport costs in the model at hand, as $\mathbf{D}$ completely contains the underlying transport cost structure.
In the case of the Beckmann model, the term $-\beta\mathbf{I}$ arises from the local congestion effects $x_{i}^{-\beta}$ and hence it is independent of $\mathbf{D}$. The term $\alpha\mathbf{D}$ reflects the proximity-based externalities and thus incorporates $\mathbf{D}$.
Similarly, for the Braid model, we can show that
$\mathbf{V}=\alpha\mathbf{D}-\mathbf{D}^{2}.$ (12)
This also implies $\omega=\alpha\Theta-\Theta^{2}$ as discussed earlier. The first term is the same agglomeration force as the Beckmann model. The second-order term $-\mathbf{D}^{2}$ captures the dispersion force stemming from competition across agents over dispersed demand.
These observations suggest a systematic way to distinguish the spatial scale of economic forces. Suppose the benefit matrix $\mathbf{V}$ is a polynomial in $\mathbf{D}$, such as $\mathbf{V}=c_{0}\mathbf{I}+c_{1}\mathbf{D}+c_{2}\mathbf{D}^{2}$. Then, we define an economic force as local if it appears as the coefficient of $\mathbf{I}$, and global if it arises from the terms involving powers of $\mathbf{D}$ [^9] A subtle point is that the distinction between local and global forces reflects modeling choices rather than universally applicable interpretations of the underlying economic mechanisms. For example, a “congestion force” may be a global force if it is formulated as crowding of regional facilities that are accessible from other regions through costly transport.
@@ 2.5   General equilibrium models
The benefit matrix approach is particularly effective for understanding general equilibrium models. For various models in the literature, the benefit matrix $\mathbf{V}$ can be cleanly represented by $\mathbf{D}$, which facilitates systematic comparison. Below, we consider a series of notable models assuming the symmetric two-region economy, and discuss them through the lens of local and global dispersion forces. Derivations are provided in Appendix F in the online appendix.
@@
The Helpman (1998) / Redding–Sturm (2008) model. The Helpman-Book1998 model considers an imperfectly competitive framework with a horizontally differentiated sector for tradable goods, and introduces the consumption of local, non-tradable goods (i.e., housing). We consider its variant proposed by Redding-Sturm-AER2008.
For illustration, we provide a slightly detailed derivation for the benefit matrix $\mathbf{V}$ for this model. The indirect utility of mobile workers is given by
$v_{i}(\bm{x})=x_{i}^{-\gamma}w_{i}^{\mu}\cdot\textsc{CMA}_{i}^{\frac{\mu}{\sigma-1}},$ (13)
@@
where $\mu\in(0,1)$ is consumers’ expenditure share on tradables and $\gamma\equiv 1-\mu$ is that on non-tradables, $\sigma>1$ is the elasticity of substitution of horizontally differentiated tradable varieties, $w_{i}$ is the wage in region $i$, and $\textsc{CMA}_{i}\equiv\sum_{j\in\mathcal{I}}x_{j}w_{j}^{1-\sigma}\phi_{ji}$ is the so-called “consumer market access.” The proximity matrix for this model is $\phi_{ij}=\tau_{ij}^{1-\sigma}$, where $\tau_{ij}\geq 1$ is the iceberg trade cost from region $i$ to $j$. Given $\bm{x}$, the wage $\bm{w}=(w_{i})$ is an endogenous object determined in general equilibrium with interregional trade.
The benefit matrix at $\bar{\bm{x}}$ reflects the structure of the indirect utility:
$\mathbf{V}$ $=-\gamma\mathbf{I}+\mu\mathbf{W}+\frac{\mu}{\sigma-1}\left(\mathbf{D}-(\sigma-1)\mathbf{D}\mathbf{W}\right).$ (14)
Here, $\mathbf{W}$ is the elasticity matrix for wages $\bm{w}$ at $\bar{\bm{x}}$, i.e., $\mathbf{W}\equiv\frac{\bar{x}}{\bar{w}}[\frac{\partial w_{i}}{\partial x_{j}}(\bar{\bm{x}})]_{i\in\mathcal{I},j\in\mathcal{I}}$ with $\bar{w}$ being the uniform level of nominal wages at $\bar{\bm{x}}$. It represents the general equilibrium response of wages to marginal migration, and given explicitly as
$\mathbf{W}=\frac{1}{\sigma}C(\mathbf{D})\cdot\mathbf{D},\quad\text{where}\quad C(\mathbf{D})\equiv\left(\mathbf{I}+\frac{\sigma-1}{\sigma}\mathbf{D}\right)^{-1}.$ (15)
In Eq. 14, the first component of $\mathbf{V}$ is a local dispersion force arising from crowding in housing markets. The second term is the response of wages. The third component is the so-called “price index effects,” the global forces driven by consumer market access. That is, the cost of living in a region falls when nearby regions offer a greater variety of goods, but rises when they pay higher wages, thereby pushing up local prices.
Substituting Eq. 15 into Eq. 14 and rearranging, $\mathbf{V}$ reduces to
$\mathbf{V}=C(\mathbf{D})\cdot\left(-\gamma\mathbf{I}+c_{1}\mathbf{D}\right),$ (16)
[…]
The final expression of $\omega$ includes a complicating term $C(\Theta)$ stemming from general equilibrium. However, since $C(\Theta)>0$, the stability of $\bar{\bm{x}}$ hinges on the sign of the linear expression $-\gamma+c_{1}\Theta$. Observe that there is a local dispersion force $-\gamma$. Also, $c_{1}$ encapsulates both positive and negative effects concerning spatial agglomeration. If $\mu$ is sufficiently large, then $c_{1}>0$ and there is a (net) global agglomeration force. [^11] In general, the sign of $c_{1}$ is indeterminate as there are both agglomeration and dispersion forces. Each term in $c_{1}$ has a clear economic interpretation as extensively discussed elsewhere (see, e.g., Fujita-Thisse-Book2013; Baldwin-et-al-Book2003; Brakman-etal-Book2019, for surveys). Since model-specific interpretations are not our main focus, such discussions are relegated to Appendix F in the online appendix. If $\sigma(1-\mu)<1$, $\bar{\bm{x}}$ is unstable for high transport costs and stable for low transport costs.
Although the key coefficients $-\gamma$ and $c_{1}$ are microfounded through general equilibrium, the implication on the stability of $\bar{\bm{x}}$ resembles that of the Beckmann model. The central dispersion mechanism in the Helpman framework is crowding in non-tradables markets, whose archetype is local congestion as in the Beckmann model.
The Allen–Arkolakis (2014) model. Allen-Arkolakis-QJE2014 proposed a perfectly competitive framework with externalities in the form of within-region congestion and positive productivity spillovers. As shown in the original paper, this model bears a structural similarity to the Helpman framework. In fact, we can show that
$\mathbf{V}=C(\mathbf{D})\cdot\left(c_{0}\mathbf{I}+c_{1}\mathbf{D}\right)$ (17)
where $C(\mathbf{D})$ captures general equilibrium adjustments, and we set $c_{0}\equiv\alpha+\beta-\frac{1+\alpha}{\sigma}$ and $c_{1}\equiv\alpha+\beta+\frac{1-\beta}{\sigma}$. Here, $\alpha>0$ and $\beta<0$ denote the elasticities of local agglomeration and congestion effects with respect to local population, respectively. As in the Helpman model, the sign of the utility gain $\omega$ again depends on a linear expression $c_{0}+c_{1}\Theta$. If agglomeration forces are sufficiently strong, the model can exhibit agglomeration under high transport costs and dispersion under low transport costs. [^12] As in the Beckmann model, sufficiently strong local dispersion force suppresses agglomeration: $|\beta|>\alpha$ implies $\omega<0$ for all transport cost levels. As in the Braid model, the symmetry is unstable for all transport cost levels if agglomeration force $\alpha$ is too large. We can thus view the Helpman model and the Allen–Arkolakis model as models driven by local dispersion forces.
The Krugman (1991) model. On the other hand, Krugman-JPE1991 emphasizes the role of market crowding, a global dispersion force, by assuming immobile consumers dispersed across regions. For this model, we compute that
$\mathbf{V}=\mathbf{W}+\frac{\mu}{\sigma-1}\left(\mathbf{D}-(\sigma-1)\mathbf{D}\mathbf{W}\right).$ (18)
Naturally, local dispersion forces are absent in $\mathbf{V}$. Furthermore, in this model,
@@
Comparing $\mathbf{W}$ in Eqs. 15 and 19, we observe that the Krugman model has a negative term $-\mathbf{D}^{2}$, while the Helpman model does not. The benefit matrix is given by

### v8

@@ 2.2  The Braid model: A “global” dispersion force
To illustrate what gain functions $\Omega$ look like under different specifications of the indirect utility function $\bm{v}$, we employ another reduced-form model. Consider replacing the local congestion term in the Beckmann model as follows:
$v_{i}(\bm{x})=y_{i}(\bm{x})\Big(\sum_{j}\phi_{ij}x_{j}\Big)^{\alpha}.$ (7)
Here, $y_{i}(\bm{x})$ is the income of an agent in the region $i$. Each region is endowed with one unit of fixed expenditure, which is allocated to all agents according to accessibility. Suppose that the share received by an agent in region $i$ from region $m$ is given by [^7] Note that the total expenditure of the region $m$ is equal to one: $\sum_{i\in\mathcal{I}}x_{i}s_{i\mid m}=1$. Possible microfoundations for this toy model are abstracted away for brevity.
$s_{i\mid m}(\bm{x})\equiv\frac{\phi_{im}}{\sum_{l}x_{l}\phi_{lm}}.$ (8)
Then, $y_{i}(\bm{x})=\sum_{m}s_{i\mid m}(\bm{x})$. We refer to the model (7) as the Braid model after Braid-RSUE1988 who studied the case $\alpha=0$.
In this model, proximity to others has a negative impact on the utility of agents, as Eq.˜8 embeds competition between agents in different regions over spatially dispersed expenditure. Concretely, for any combination of $i,j,m\in\{1,2\}$,
$\frac{\partial s_{i\mid m}}{\partial x_{j}}=-\frac{\phi_{im}}{\sum_{l}x_{l}\phi_{lm}}\cdot\frac{\phi_{jm}}{\sum_{l}x_{l}\phi_{lm}}=-s_{i\mid m}\cdot s_{j\mid m}<0,$ (9)
demonstrating that a marginal increase in agents in any region $j\in\{1,2\}$ has negative impacts on $s_{i\mid m}$. This is simply because it increases the denominator of $s_{i\mid m}$. Whether income $y_{i}=\sum_{j}s_{i\mid j}$ as a whole increases or not after a migration shock depends on the relative magnitudes of these impacts. Specifically, analogous to the utility gain $\omega$, we can compute the income gain of a marginal mover from region $2$ to 11: [^8] From Eq. 9, $\frac{\partial{y_{1}}}{\partial{x_{1}}}=-(s_{1\mid 1})^{2}-(s_{1\mid 2})^{2}=-\frac{1+\phi^{2}}{\bar{x}^{2}(1+\phi)^{2}}$ and $\frac{\partial{y_{2}}}{\partial{x_{1}}}=-s_{2\mid 1}\cdot s_{1\mid 1}-s_{2\mid 2}\cdot s_{1\mid 2}=-\frac{2\phi}{\bar{x}^{2}(1+\phi)^{2}}$.
$\frac{\bar{x}}{\bar{y}}\left.\left(\frac{\partial{y_{1}(\bm{x})}}{\partial{x_{1}}}-\frac{\partial{y_{2}(\bm{x})}}{\partial{x_{1}}}\right)\right|_{\bm{x}=\bar{\bm{x}}}=-\frac{1+\phi^{2}}{(1+\phi)^{2}}+\frac{2\phi}{(1+\phi)^{2}}=-\frac{(1-\phi)^{2}}{(1+\phi)^{2}}=-\Theta^{2}<0,$ (10)
where $\bar{y}=y_{1}(\bar{\bm{x}})=y_{2}(\bar{\bm{x}})=1/\bar{x}$. This shows that a marginal migration shock always induces an income loss for the movers. This force is on the second order of $\Theta$, reflecting that the proximity matrix $\phi_{ij}]$ appears twice in the numerator of Eq.˜9.
From Eqs.˜4 and 10, the utility gain for the Braid model is given as follows:
@@
The gain function for the model is therefore $\Omega(\Theta)=\alpha\Theta-\Theta^{2}$. The first term is the same agglomeration force as in the Beckmann model. The second term, $\Theta^{2}$, represents the “global” dispersion force due to the income loss in Eq.˜10.
This dispersion force weakens as $\Theta$ falls (i.e., as $\phi$ increases). If $\phi$ is very small, regions are effectively in autarky and $y_{i}\approx s_{i\mid i}\approx 1/x_{i}$. In this case, a migration shock that raises $x_{i}$ directly intensifies local competition and lowers income, creating a strong incentive to disperse ($\Theta^{2}\approx-1$). By contrast, if $\phi$ is close to one, regions face nearly identical crowding conditions ($y_{i}\approx 1$), and migration shocks have little effect on income ($\Theta^{2}\approx 0$).
In the model, $\bar{\bm{x}}$ can become unstable for some $\phi$ whenever $\alpha>0$. If $0<\alpha<1$, $\bar{\bm{x}}$ is stable for high transport costs (large $\Theta$ $\Leftrightarrow$ small $\phi$) and unstable for low transport costs (small $\Theta$ $\Leftrightarrow$ large $\phi$). If $\alpha\geq 1$, $\bar{\bm{x}}$ is unstable for all $\phi$.
Contrasting implications of transportation costs. The two reduced-form models yield fundamentally opposing implications. In the Beckmann model, $\bar{\bm{x}}$ is stable for low transport costs (large $\phi$), and agglomeration occurs for high transport costs (small $\phi$). The Braid model exhibits the opposite behavior. Figures˜4B and 4A confirm this by showing the full equilibrium paths on the $\phi$-axis. The contrast persists in the presence of asymmetries. For instance, if region 1 possesses an exogenous advantage as in Figs.˜4D and 4C, a symmetric configuration is no longer an equilibrium. Nevertheless, the qualitative findings remain consistent with the symmetric case: an increase in $\phi$ (declining transport costs) fosters dispersion in the Beckmann model but drives agglomeration in the Braid model.
This divergence is rooted in the different spatial scales of dispersion forces. In the Beckmann model, dispersion under high $\phi$ is driven by the relative weakening of the agglomeration force, whereas the “local” dispersion force remains invariant to $\phi$. In the Braid model, endogenous agglomeration occurs due to the relative decline of the “global” dispersion force as $\phi$ increases. Thus, the spatial scale of the dominant dispersion force can alter the implications of declining transport costs. [^9] In fact, the literature on two-region models has recognized that various dispersion mechanisms can have opposing effects. For example, Fujita-Thisse-Book2013 compares the seminal models by Krugman-JPE1991 and Helpman-Book1998 and noted that “Krugman’s scenario is reversed” (p.289) in Helpman-type frameworks with urban costs. As noted in the introduction and further discussed in Section˜3, in many-region settings, these two types of dispersion forces lead to contrasting spatial patterns.
@@ D The Braid model with asymmetry
Figure 4: Equilibrium values of $x_{1}$ in the Beckmann model and the Braid model
@@ 2.3  Benefit matrix and the spatial scale of economic forces
While the insights from the stylized models considered so far are intuitive and nearly immediate, they rely on a reduced-form structure. In more comprehensive frameworks, particularly general equilibrium models where wages, prices, and land rents are endogenously determined, the impact of transport costs becomes considerably more complex. In such settings, multiple forces interact simultaneously, making it difficult to isolate whether “local” or “global” dispersion dominates.
To bridge the gap between our stylized insights and richer general equilibrium specifications, it is necessary to develop a formal method for evaluating how these endogenous forces respond to changes in transport costs. For this purpose, we introduce an analytical tool termed the benefit matrix. This matrix allows us to systematically decompose the spatial externalities inherent in spatial models and classify them into the local and global forces in our earlier discussion.
Definition 1 (Benefit matrix).
For a differentiable indirect utility function $\bm{v}$, its benefit matrix $\mathbf{V}$ is its elasticity matrix at symmetric equilibrium $\mathbf{V}\equiv\frac{\bar{x}}{\bar{v}}[\frac{\partial v_{i}}{\partial x_{j}}(\bar{\bm{x}})]_{i=1,2;j=1,2}$.
@@
Then, the utility gain $\omega$ in the symmetric equilibrium $\bar{\bm{x}}$ is the eigenvalue of the benefit matrix $\mathbf{V}$, associated with the eigenvector $\bm{z}=(1,-1)$. Intuitively, $\bm{z}$ represents the direction of possible migration shocks because $(\bar{x}+\epsilon,\bar{x}-\epsilon)=\bar{\bm{x}}+\epsilon\bm{z}$.
Example 1.
Consider the Beckmann model. Its benefit matrix is
$\mathbf{V}=-\beta\mathbf{I}+\alpha\mathbf{D},$ (12)
where $\mathbf{I}$ is the identity matrix and $\mathbf{D}$ is the row-normalized proximity matrix:
$\mathbf{D}=\frac{1}{1+\phi}\begin{bmatrix}1&\phi\\ \phi&1\end{bmatrix}$. (13)
@@
We confirm $\mathbf{D}\bm{z}=\frac{1-\phi}{1+\phi}\bm{z}=\Theta\bm{z}$. That is, $\bm{z}=(1,-1)$ is an eigenvector of $\mathbf{D}$, and the proximity gain $\Theta$ is the associated eigenvalue. Then, $\mathbf{V}\bm{z}=(-\beta+\alpha\Theta)\bm{z}=\omega\bm{z}$ from Eq.˜12, showing that $\omega$ is indeed an eigenvalue of $\mathbf{V}$ corresponding to $\bm{z}$. ∎
Example 2.
The benefit matrix for the Braid model is
$\mathbf{V}=\alpha\mathbf{D}-\mathbf{D}^{2}.,$ (14)
Since $\mathbf{V}\bm{z}=(\alpha\Theta-\Theta^{2})\bm{z}$, $\omega=\alpha\Theta-\Theta^{2}$ is the eigenvalue of $\mathbf{V}$ associated with $\bm{z}$. ∎
Examples˜1 and 2 illustrate that the gain function $\Omega$ introduced earlier (Eq.˜6) arises from the structure of $\mathbf{V}$. The utility gain takes the form $\omega=\Omega(\Theta)$ because the benefit matrix itself can be written in parallel form $\mathbf{V}=\Omega(\mathbf{D})$. Here, the matrix function is interpreted in a straightforward way, e.g., matrix polynomials. We can thus think of $\mathbf{V}=\Omega(\mathbf{D})$ and $\omega=\Omega(\Theta)$ interchangeably.
In the two reduced-form examples, the gain function $\Omega(\Theta)$ allows us to formally distinguish the spatial scale of the economic forces. A negative constant term in the gain function $\Omega(\Theta)$ corresponds to a local dispersion force. In contrast, negative non-constant terms correspond to global dispersion forces. The spatial scale for agglomeration forces can be similarly defined: positive constants are local agglomeration forces, and negative non-constant terms are global agglomeration forces.
@@ 2.4  General equilibrium models
We now examine specific models in the literature to illustrate that their benefit matrices $\mathbf{V}$ are simple functions of $\mathbf{D}$. We can then immediately obtain the associated gain functions and systematically determine the spatial scale of the dispersion forces in each model. Detailed derivations are provided in Online Appendix F.
@@
The Helpman-Book1998 / Redding-Sturm-AER2008 model. Helpman considered an imperfectly competitive framework in which agents consume both differentiated tradable goods and local non-tradable goods (i.e., housing). We consider its variant by Redding and Sturm. The indirect utility of mobile workers in this model is
$v_{i}(\bm{x})=x_{i}^{-(1-\mu)}w_{i}^{\mu}\cdot\textsc{CMA}_{i}^{\frac{\mu}{\sigma-1}},$ (15)
@@
where $\mu\in(0,1)$ is consumers’ expenditure share on tradables and $1-\mu\in(0,1)$ is that on non-tradables, $\sigma>1$ is the elasticity of substitution of horizontally differentiated tradable varieties, $w_{i}$ is the nominal wage in region $i$, and $\textsc{CMA}_{i}\equiv\sum_{j\in\mathcal{I}}x_{j}w_{j}^{1-\sigma}\phi_{ji}$ is the so-called “consumer market access.” The proximity matrix is $\phi_{ij}=\tau_{ij}^{1-\sigma}$, where $\tau_{ij}\geq 1$ is the iceberg trade cost from region $i$ to $j$. Given $\bm{x}$, the wage $\bm{w}=(w_{i})$ is endogenously determined in general equilibrium with interregional trade.
The benefit matrix for this model can be computed as follows:
$\mathbf{V}=C(\mathbf{D})\cdot\left(-(1-\mu)\mathbf{I}+c_{1}\mathbf{D}\right),$ (16)
where $C(\mathbf{D})\equiv\big(\mathbf{I}+\frac{\sigma-1}{\sigma}\mathbf{D}\big)^{-1}$ and $c_{1}\equiv\frac{\mu}{\sigma-1}+\frac{\mu}{\sigma}-(1-\mu)\frac{\sigma-1}{\sigma}$. This then implies
$\omega=C(\Theta)\cdot\left(-(1-\mu)+c_{1}\Theta\right)$ (17)
where $C(\Theta)\equiv(1+\frac{\sigma-1}{\sigma}\Theta)^{-1}>0$. Since $C(\Theta)>0$, the sign of $\omega$ hinges on the sign of
$\omega^{\sharp}\equiv-(1-\mu)+c_{1}\Theta,$ (18)
which is similar to the Beckmann model. In particular, $(1-\mu)$ corresponds to the local dispersion force due to crowding in the non-tradables market in each region. If $\mu$ is sufficiently large, then agents’ love for variety produces a strong global agglomeration force: we have $c_{1}>0$ and hence $c_{1}\Theta>0$ if $\mu>(\frac{\sigma-1}{\sigma})^{2}$, where we note $\frac{\sigma-1}{\sigma}\in(0,1)$. If further $\mu>\frac{\sigma-1}{\sigma}$, $\bar{\bm{x}}$ is unstable for high transport costs (large $\Theta$) and stable for low transport costs (small $\Theta$), just as in the Beckmann model.
We provide a slightly detailed derivation behind the final expressions (16) and (17) for illustration. The definition of $\bm{v}$ in Eq.˜15 yields
$\mathbf{V}$ $=-(1-\mu)\mathbf{I}+\mu\mathbf{W}+\frac{\mu}{\sigma-1}\left(\mathbf{D}-(\sigma-1)\mathbf{D}\mathbf{W}\right).$ (19)
The first term in Eq.˜19 represents the crowding in the housing markets. The second is the direct impact of nominal income on indirect utility, where
$\mathbf{W}=\frac{1}{\sigma}C(\mathbf{D})\cdot\mathbf{D}$ (20)
is the elasticity matrix of nominal wages. [^10] Concretely, $\mathbf{W}\equiv\frac{\bar{x}}{\bar{w}}[\frac{\partial w_{i}}{\partial x_{j}}(\bar{\bm{x}})]_{i=1,2;j=1,2}$ with $\bar{w}=w_{1}(\bar{\bm{x}})=w_{2}(\bar{\bm{x}})$. As $C(\mathbf{D})=\big(\mathbf{I}+\frac{\sigma-1}{\sigma}\mathbf{D}\big)^{-1}$ plays a role analogous to the Leontief inverse in input–output analysis, $\mathbf{W}$ captures the general equilibrium response of wages to marginal migration shocks taking into account interregional trade. The third term in Eq.˜19 is the “cost-of-living” effects: the price index in a region falls when nearby regions offer a greater variety of goods but increases when they pay higher wages and thus charge higher prices (see, e.g., Fujita-Thisse-Book2013; Baldwin-et-al-Book2003; Brakman-etal-Book2019).
Substituting Eq.˜20 into Eq.˜19 and rearranging, we obtain
$\mathbf{V}$ $=-(1-\mu)\mathbf{I}+\frac{\mu}{\sigma-1}\mathbf{D}+\frac{\mu}{\sigma}(\mathbf{I}-\mathbf{D})\cdot C(\mathbf{D})\cdot\mathbf{D}\quad\text{and hence}$ (21)
$\omega$ $=\Omega(\Theta)=\underbrace{-(1-\mu)}_{<0}\,+\,\underbrace{\frac{\mu}{\sigma-1}\Theta}_{>0}\,+\,\underbrace{\frac{\mu}{\sigma}C(\Theta)(1-\Theta)\Theta}_{>0}.$ (22)
The first two terms represent the partial equilibrium utility gains where wage adjustments are ignored. The third term in Eq.˜22 summarizes the net impact of wages on indirect utility in general equilibrium, both through the (positive) individual-level income gain and the (negative) cost-of-living effect. It is a net global agglomeration force because it is strictly positive for all $\Theta\in(0,1)$. Thus, the only dispersion force that can stabilize $\bar{\bm{x}}$ is the local dispersion force captured by the first term of Eq.˜22.
Further rearrangement of Eqs.˜21 and 22 yields the final expressions (16) and (17). By construction, $\omega^{\sharp}$ in Eq.˜18 captures the net utility gain considering all endogenous forces and their general equilibrium trade-offs. For example, despite there is no net global dispersion force, $\omega^{\sharp}=-(1-\mu)+c_{1}\Theta$ has a negative term involving $\Theta$:
$c_{1}\Theta=\frac{\mu}{\sigma}\Theta+\frac{\mu}{\sigma-1}\Theta\underbrace{-(1-\mu)\frac{\sigma-1}{\sigma}\Theta}_{<0}$ (23)
This last term represents how the local dispersion force $(1-\mu)$ counteract, in each agent’s migration incentives, the core agglomeration forces in the model, such as the love for variety and demand linkage (the first two terms of Eq.˜23).
The Allen-Arkolakis-QJE2014 model. The model is a perfectly competitive framework with both positive and negative externalities. Appendix F.2.4 shows that the benefit matrix for the Allen–Arkolakis model is
$\mathbf{V}=C(\mathbf{D})\cdot\left(c_{0}\mathbf{I}+c_{1}\mathbf{D}\right),$ (24)
where $C(\mathbf{D})\equiv((\sigma\mathbf{I}+(\sigma-1)\mathbf{D})(\mathbf{I}-\mathbf{D}))^{-1}$, $c_{0}\equiv\alpha-\beta-\frac{1+\alpha}{\sigma}$, and $c_{1}\equiv\alpha-\beta+\frac{1+\beta}{\sigma}$. The parameters $\alpha>0$ and $\beta>0$ are the magnitudes of local agglomeration and congestion effects with respect to local population, respectively, and $\sigma>1$ is the elasticity of substitution across regionally differentiated goods. The gain function $\Omega$ is obtained accordingly. Analogous to the Helpman model, the sign of $\omega$ depends on a linear expression $\omega^{\sharp}\equiv c_{0}+c_{1}\Theta$ that captures the net migration incentive for agents.
If the dispersion force is relatively strong (relatively large $\beta$), we have $c_{0}<0$, indicating a net local dispersion force. Likewise, if the agglomeration force is relatively strong (relatively large $\alpha$), we have $c_{1}>0$, indicating a net global agglomeration force. In particular, if $\beta<\alpha<\bar{\alpha}\equiv\frac{\beta\sigma+1}{\sigma-1}$, agglomeration occurs if transport costs are high (small $\phi$) and dispersion occurs if transport costs are low (large $\phi$). Under the presence of externalities, the only stabilizing force in the Allen–Arkolakis model is the local crowding effect. [^11] In the perfectly competitive case ($\alpha=0$ and $\beta=0$), the model reduces to the Armington-IMF1969 framework. For this case, $\omega=\frac{1}{\sigma}C(\Theta)(-1+\Theta)<0$, and the net agglomeration term $(-1+\Theta)$ does not contain any negative components in $\Theta$. Nonetheless, $\omega$ is a negative function and can be interpreted as the underlying “global dispersion force” inherent to general equilibrium in the Armington framework. That is, as discussed in Allen-Arkolakis-QJE2014, this model bears a structural similarity to the Helpman model.
The Krugman-JPE1991 model. Krugman’s seminal model emphasizes the role of market crowding. In this model, there is always a nonzero demand for the manufacturing goods in each region due to the presence of immobile consumers. This discourages concentration of production in one region if transport costs are high, and this produces a global dispersion force as in the Braid model. For this model, the benefit matrix is
$\mathbf{V}=\mathbf{W}+\frac{\mu}{\sigma-1}\left(\mathbf{D}-(\sigma-1)\mathbf{D}\mathbf{W}\right)$ (25)
where the elasticity matrix for the nominal wage is
@@
Compared to Eq.˜19, the local dispersion force is absent. Compared with Eqs.˜20 and 26, the Krugman model has a negative term $\mathbf{D}^{2}$, while the Helpman model does not. We can rearrange $\mathbf{V}$ to see